# Language Models Know the Value of Numbers

### v3

@@ Abstract
Large language models (LLMs) have exhibited impressive competence in various tasks, but their internal mechanisms on mathematical problems are still under-explored. In this paper, we study a fundamental question: whether language models know the value of numbers, a basic element in math. To study the question, we construct a synthetic dataset comprising addition problems and utilize linear probes to read out input numbers from the hidden states. Experimental results support the existence of encoded number values in LLMs on different layers, and these values can be extracted via linear probes. Further experiments show that LLMs store their calculation results in a similar manner, and we can intervene the output via simple vector additions, proving the causal connection between encoded numbers and language model outputs. Our research provides evidence that LLMs know the value of numbers, thus offering insights for better exploring, designing, and utilizing numeric information in LLMs.
Language Models Know the Value of Numbers
@@ 1 Introduction
Numbers are fundamental elements in math, and the way how LLMs process numbers essentially influences the final outcome. In order to accurately answer mathematical problems, LLMs need the ability to get the precise value of numbers in the input text. Currently, the way how LLMs process numbers is still not fully explored. While previous studies Stolfo et al. (2023); Hanna et al. (2024) have explored the inner mechanisms of language models on mathematical problems, they focus on small numbers that can be represented with a single token, and how LLMs utilize multi-token large numbers remains largely unknown.
@@
In this paper, we explore the question whether LLMs know the value of numbers through extracting numerical information from their internal representations. To be specific, we construct a synthetic dataset comprising simple addition questions, and train linear probes Nanda et al. (2023); Gurnee and Tegmark (2023) on the hidden states of LLMs to predict the number values provided in the input text. Experimental results on the dataset demonstrate that the value of input numbers can be probed from the hidden states of language models from early layers. Both input values and calculation results can be read out, and encoded values can be found at different positions in the input token sequence. These experimental results support that language models do encode numerical information.
@@
To sum up, our contributions can be listed as: (1) We study the question of whether language models are able to encode the value of numbers in the input text and construct a synthetic dataset to analyze the language models. (2) We discover that language models do encode the value of numbers by utilizing linear probes to probe encoded number values in the hidden states of language models. (3) We further prove that language models utilize the encoded numerical information by revealing the causal connection between encoded number values and the final output of language models.
@@ 2.1 The Goal of Probing
where $f_{i}$ refers to the process of encoding on layer $i$, and $t-x$ refers to the remaining part in $t$ aside from $x$.
@@ 2.2 Dataset Construction
To investigate whether LLMs know numbers, we construct a synthetic dataset containing different magnitudes of numbers. The dataset contains numbers ranging from 2 digits to 10 digits, with each digit corresponding to 1000 entries [^1] See Appendix A for more details.. We split the dataset into training, validation, and test sets at a ratio of 80%/10%/10%.
To observe how LLMs encode and utilize numbers, we adopt addition problems as our prompt [^2] We do not include subtraction problems in our dataset in that subtraction problems are similar to addition problems. We do not include multiplication and division problems either, as LLMs perform poorly on these problems (even 5-digit multiplication yields an accuracy of about 0%.. Let $a$ and $b$ be two randomly generated numbers, each question is formulated as:
@@ 2.5 Experimental Setup
In main experiments, we probe 3 distinct values at different positions: the first number $a$ at the number token of $a$, the second number $b$ at the last number token of $b$, and the prediction of language models $o$ at the last token of the entire input text. We report the accuracy of $o$, i.e. the ratio of $o=a+b$, in Appendix C.
@@
3 Do LLMs Know Number Values?
@@ 3.1 The Existence of Encoded Number Values
This phenomenon may indicate that language models use stronger non-linear encoding systems, which we will further explore in Section 3.3; Or it may be a hint that the number encoding in language models is not precise [^4] See Appendix E for more detailed experiments..
@@
Figure 4: How the pattern of probe predictions on the first input number $a$ changes as the layer gets deeper.
@@ 3.2 Number Encoding Patterns are Different across Layers
Figure 5: The comparison between linear probes and MLP probes on mean square error (MSE). The MLP probes do not show advantage over linear probes. More detailed experiments are reported in Appendix D.
@@ 3.3 LLMs Encode Numbers Linearly
3.4 Numeric Information Persist at Subsequent Positions
(a) MSE of probes on $a$.
(b) MSE of of probes on $b$.
@@ 4.1 Patching Encoded Numbers
We adopt the activation patching technique proposed by Stolfo et al. (2023) to quantify the importance of encoded number values $\mathbf{h}_{i}$ at different layers $i$ and different token positions.
@@
which measures how much a specific layer $i$ at position $t$ contributes to the final shifting of results.
@@
Language models concern only certain tokens during calculation.
@@
Each position can construct the number encoding from scratch.
The effect of patching on different number digits displays a clear pattern: the earlier a digit appears, the higher the patching effect is. A possible explanation is that each position constructs its number encoding from scratch, and their encoded number values are independent from each other. For example, although the activation at digit “3” in “1234” encodes the value of 123, language models will only attend to the value 2 in the encoding process at subsequent positions, which is demonstrated in Table 1. More detailed experiments are reported in Appendix G.
@@
The effect of activation patching is clearly divided into two parts: in early layers before layer 14, patching the number tokens greatly influences the final outcome, while patching the last token is mostly ineffective; but in late layers after layer 20 it is just the opposite. We can assume that early layers perform the task of processing the value of input number token sequences, while late layers use encoded values to calculate the final outcome, which is similar to the findings in Stolfo et al. (2023).
Figure 8: The success rate of performing a linear intervention on 6 consecutive layers. The red line represents the results of the baseline model which simply extends the hidden states in their original directions. More detailed experiments are reported in Appendix H.
@@ 4.2 Linearly Intervening Encoded Numbers
To determine whether the encoded computational results causally affect the outcome of language models, we conduct linear intervention experiments similar to Nanda et al. (2023). For each intervened layer $i$, we add the number encoding direction vector $\mathbf{d}_{i}$ to the residual stream $\mathbf{h}_{i}$:
@@
We also record the results of taking normalized $\mathbf{h}_{i}$ as the intervention direction $\mathbf{d}_{i}$, which represents a null intervention method for comparison.
@@ 5 Discussion and Future Directions
It seems that LLMs are currently not able to precisely encode numbers, and the compression loss will inevitably bring errors to subsequent computation, especially when the input numbers are large. Developing specialized encoding systems that could give precise presentations for numbers Golkar et al. (2023) could eliminate errors at the root, thus helping LLMs better solve mathematical problems.
@@
By adding modules that directly utilize the encoded numbers in language models, the computational errors may be further reduced, especially on large-number calculations. We conduct a pioneer experiment in Appendix I to reveal the potential of controlling computational errors with probed numbers.
@@ 6 Related Work
How language models process numbers has been studied by multiple researchers. Wallace et al. (2019) detects the existence of numeracy in static pre-trained word embeddings. Hanna et al. (2024) finds a critical circuit that performs greater-than comparisions in GPT-2. Stolfo et al. (2023) studies how language models process arithmetic information by intervening on specific modules of the model.
@@ 7 Conclusion
In this paper, we study the question of whether large language models know the value of numbers. If number values can be extracted from the internal representations of LLMs, we can assume that language models to some extent know the value of numbers. We construct a dataset consisting of simple addition problems and introduce linear probes to investigate whether language models encode number values.
Experimental results prove that LLMs do encode the value of input numbers, and the representation could be linearly read out. The ability to encode numbers is consistent across different model scales, and the encoding seems to be the most precise on intermediate layers. Further experiments show that LLMs utilize the encoded number values to perform arithmetic calculations, and the behavior of language models can be controlled via simple linear interventions, proving the causal connection between encoded numbers and model outputs.
@@ Appendix B Experiment Implementation
Appendix C Overall Accuracy
@@
Appendix D Detailed Experiments on Linearity
@@
Appendix E Experimental Results on Partial Number Encoding
@@
Appendix F Probing With Control Tasks
@@
Appendix G Detailed Experiments on Activation Patching
G.1 Examples of Patching
@@
Appendix H Detailed Experiments on Linear Intervention
@@ 25-30 Answer: 6455
We also observe that despite intervening on early or late layers both lead to poor success rates, and they display different patterns of output. Table 3 shows the result of intervening on different layers of Mistral-7B. It can be seen that performing a linear intervention on early layers would completely destroy the final outcome, while intervening on late layers will not change the result at all. We hypothesize that the number encoding in early layers has not fully developed yet, and intervening in it would lead to unexpected results; In late layers, the number encoding is simply remembered but not used, and the language models rely on other subspace to decode the final outcome.
@@
We observe that on layer 14 of Mistral-7B, the L2 norm of $\mathbf{h}_{i}$ is around 4, while the L2 norm of $\mathbf{W}_{i}$ is around 10. Thus, the norm of $\mathbf{d}_{i}$ would be around 0.1, which is way smaller than the norm of $\mathbf{h}_{i}$. A possible hypothesis is that the probed number value is the projection of $\mathbf{h}_{i}$ along the direction $\mathbf{W}_{i}$, and simply adding vectors to $\mathbf{h}_{i}$ would draw it away from its valid subspace. To maintain intervened $\mathbf{h}_{i}$ in its valid subspace, it should be rotated along some direction. The method of precisely changing the encoded number values in language models still remains to be explored.
@@
Appendix I Directly Calculate with Encoded Number Values

### v4

# Language Models Encode the Value of Numbers Linearly

@@ Abstract
Large language models (LLMs) have exhibited impressive competence in various tasks, but their internal mechanisms on mathematical problems are still under-explored. In this paper, we study a fundamental question: how language models encode the value of numbers, a basic element in math. To study the question, we construct a synthetic dataset comprising addition problems and utilize linear probes to read out input numbers from the hidden states. Experimental results support the existence of encoded number values in LLMs on different layers, and these values can be extracted via linear probes. Further experiments show that LLMs store their calculation results in a similar manner, and we can intervene the output via simple vector additions, proving the causal connection between encoded numbers and language model outputs. Our research provides evidence that LLMs encode the value of numbers linearly, offering insights for better exploring, designing, and utilizing numeric information in LLMs. The code and data are available at https://github.com/solitaryzero/NumProbe.
Language Models Encode the Value of Numbers Linearly
@@ 1 Introduction
Numbers are fundamental elements in math. In order to accurately answer mathematical problems, LLMs should be able to precisely encode value of numbers in the input text. Currently, the way how LLMs process numbers is still not fully explored. While previous studies Stolfo et al. (2023); Hanna et al. (2024) have explored the inner mechanisms of language models on mathematical problems, they focus on small numbers or modular arithmetic Engels et al. (2024); Zhong et al. (2024), and how LLMs utilize numbers in a larger, unconstrained range remains largely unknown.
@@
In this paper, we explore the question whether and how LLMs encode the value of numbers through extracting numerical information from their internal representations. To be specific, we construct a synthetic dataset comprising simple addition questions, and train linear probes Nanda et al. (2023); Gurnee and Tegmark (2023) on the hidden states of LLMs to predict the number values provided in the input text. Experimental results on the dataset demonstrate that the value of input numbers can be probed from the hidden states of language models from early layers, as illustrated in Figure 1. Both input values and calculation results can be read out, and encoded values can be found at different token positions. These results support that language models do encode numerical information, and possibly in a linear manner.
@@
To sum up, our contributions can be listed as: (1) We study the question of whether language models are able to encode the value of numbers in the input text and construct a synthetic dataset to analyze the language models. (2) We discover that language models encode the value of numbers linearly by utilizing linear probes to probe encoded number values in the hidden states of language models. (3) We further prove that language models utilize the encoded numerical information by revealing the causal connection between encoded number values and the final output of language models.
@@ 2.1 The Goal of Probing
where $f_{i}$ refers to the encoding process on layer $i$, and $t-x$ refers to the tokens in $t$ excluding $x$.
@@ 2.2 Dataset Construction
To investigate whether LLMs encode numbers, we construct a synthetic dataset containing different magnitudes of numbers. The dataset contains numbers ranging from 2 digits to 10 digits, with each digit corresponding to 1000 entries [^1] See Appendix A for more details.. We split the dataset into training, validation, and test sets at a ratio of 80%/10%/10%.
To observe how LLMs encode and utilize numbers, we adopt addition problems as our prompt [^2] The experimental results on subtraction problems are similar to the results on addition problems (see Appendix C). We do not include multiplication and division problems either, as LLMs perform poorly on these problems (even 5-digit multiplication yields an accuracy of about 0%).. Let $a$ and $b$ be two randomly generated numbers, each question is formulated as:
@@ 2.5 Experimental Setup
In main experiments, we probe 3 distinct values at different positions: the first number $a$ at the last digit of $a$ (for example, 3 for 123), the second number $b$ at the last digit of $b$, and the prediction of language models $o$ at the last token of the entire input text. We report the accuracy of $o$, i.e. the ratio of $o=a+b$, in Appendix D.
@@
3 Do LLMs Encode Number Values?
@@ 3.1 The Existence of Encoded Number Values
This phenomenon may indicate that language models use stronger non-linear encoding systems, which we will further explore in Section 3.4; Or it may be a hint that the number encoding in language models is not precise [^4] See Appendix F for more detailed experiments..
@@
Figure 4: How the pattern of probe predictions on the first input number $a$ changes as the layer gets deeper. Probe predictions on different layers of LLaMA-2-7B show different patterns.
(a) MSE of probes on $a$.
(b) MSE of of probes on $b$.
@@ 3.2 Number Encoding Patterns are Different across Layers
3.3 Numeric Information Persist at Subsequent Positions
Figure 6: The comparison between linear probes and MLP probes on mean square error (MSE). The MLP probes do not show advantage over linear probes. More detailed experiments are reported in Appendix E.
@@ 4.1 Patching Encoded Numbers
Firstly, we study the influence of number encoding at different positions by changing the activation of language models. We adopt the activation patching technique proposed by Stolfo et al. (2023) to quantify the importance of encoded number values $\mathbf{h}_{i}$ at different layers $i$ and different token positions.
@@
which measures how much a specific layer $i$ at position $t$ affects the final result. Note that the metric is intended for qualitative rather than quantitative analysis.
@@
Each digit affects the result independently.
The effect of patching on different number digits displays a clear pattern: the earlier a digit appears, the more patching it changes the final output value. While the latter digits encode the values of partial number sequences (See Appendix F for details), activation patching seems to only change the final result by the value of the digit itself. For example, although the activation at digit “3” in “1234” encodes the value of 123, patching it equals changing the input number to "1294" rather than "9994", as demonstrated in Table 1. More detailed experiments are reported in Appendix H.
Language models concern only certain tokens during calculation.
@@
The effect of activation patching can be divided into two parts: on early layers before layer 14, patching the number tokens greatly influences the final outcome, while patching the last token is mostly ineffective; but in late layers after layer 20 it is just the opposite. We assume that early layers perform the task of processing the value of input number token sequences, while late layers use encoded values to calculate the final outcome, which is similar to the findings in Stolfo et al. (2023).
Figure 8: The success rate of performing a linear intervention on 6 consecutive layers. More detailed experiments are reported in Appendix I.
@@ 4.2 Linearly Intervening Encoded Numbers
To determine whether the encoded computational results causally affect the outcome of language models, we linearly intervene the hidden states and see whether the output changes as expected.
@@
Following the method of Nanda et al. (2023), for each intervened layer $i$, we add the number encoding direction vector $\mathbf{d}_{i}$ to the residual stream $\mathbf{h}_{i}$:
@@
We also use two alternative directions as baselines: normalized $\mathbf{h}_{i}$ as null intervention, and a random unit vector $\mathbf{I}$ as random intervention.
@@ 5 Discussion and Future Directions
The loss of encoded number values in LLMs will inevitably bring errors to subsequent computation, especially for large input numbers. Developing specialized encoding systems that could give precise presentations for numbers Golkar et al. (2023) could eliminate errors at the root, thus helping LLMs better solve mathematical problems.
@@
By adding modules that directly utilize the encoded numbers in language models, the computational errors may be further reduced, especially on large-number calculations. We conduct a pioneer experiment in Appendix J to reveal the potential of controlling computational errors with probed numbers.
@@ 6 Related Work
How language models process numbers has been studied by multiple researchers. Wallace et al. (2019) detects the existence of numeracy in static pre-trained word embeddings. Hanna et al. (2024) finds a critical circuit that performs greater-than comparisions in GPT-2. Stolfo et al. (2023) studies how language models process arithmetic information by intervening on specific modules of the model. Zhong et al. (2024); Engels et al. (2024) discover evidence that numbers on modular arithmetic may be circularly encoded.
@@ 7 Conclusion
In this paper, we study the question of whether and how large language models encode the value of numbers. If number values can be extracted from the internal representations of LLMs, we can assume that LLMs encode the value of numbers in their hidden states. We construct a dataset consisting of simple addition problems and introduce linear probes to investigate whether language models encode number values.
Experimental results prove that LLMs do encode the value of input numbers, and the representation could be linearly read out. The ability to linearly encode numbers is consistent across different model scales, and the encoding seems to be the most precise on intermediate layers. Further experiments show that LLMs utilize the encoded number values to perform arithmetic calculations, and the behavior of language models can be controlled via simple linear interventions, proving the causal connection between encoded numbers and model outputs.
@@ Appendix B Experiment Implementation
(a) $\rho$ of probes on $a$.
(b) $\rho$ of probes on $b$.
(c) $\rho$ of probes on $o$.
(d) $R^{2}$ of probes on $a$.
(e) $R^{2}$ of probes on $b$.
(f) $R^{2}$ of probes on $o$.
(g) AAcc of probes on $a$.
(h) AAcc of probes on $b$.
(i) AAcc of probes on $o$.
(j) MSE of probes on $a$.
(k) MSE of probes on $b$.
(l) MSE of probes on $o$.
Figure 10: Pearson coefficient ($\rho$), out-of-sample $R^{2}$, approximate accuracy (AAcc), and mean square error (MSE) of probes on different layers for subtraction problems. $a$ and $b$ refer to the two input numbers denoted in Section 2.2, and $o$ refers to the prediction of language models respectively. High $\rho$ and $R^{2}$ indicate the existence of encoded number values in the hidden states.
Appendix C Experiments on Subtraction Problems
In the main paper, we only show the results of probing on addition problems. We also conduct experiments on subtraction problems with the form of:
Question: What is the result of {a}
minus {b}?
Answer: {a - b}
where we assert $a>b$ to ensure the result being a positive number.
Figure 10 demonstrates the result of probing on subtraction problems. We can clearly observe that the trends of different metrics are similar to those on addition problems. In other words, the behaviour of language models on subtraction problems are similar to the behaviour on addition problems.
Appendix D Overall Accuracy
@@
Appendix E Detailed Experiments on Linearity
@@
Appendix F Experimental Results on Partial Number Encoding
@@
Appendix G Probing With Control Tasks
@@
Appendix H Detailed Experiments on Activation Patching
@@
Appendix I Detailed Experiments on Linear Intervention
I.1 Success Rate
@@ 25-30 Answer: 6455
I.2 Output Patterns
We also observe that while intervening on early or late layers both lead to poor success rates, they display different patterns of output. Table 3 shows the result of intervening on different layers of Mistral-7B. It can be seen that performing a linear intervention on early layers would completely destroy the final outcome, while intervening on late layers will not change the result at all. We hypothesize that the number encoding in early layers has not fully developed yet, and intervening in it would lead to unexpected results; In late layers, the number encoding is simply remembered but not used, and the language models rely on other subspace to decode the final outcome.
I.3 Additional Experiments
@@
A possible hypothesis is that the probed number value is the projection of $\mathbf{h}_{i}$ along the direction $\mathbf{W}_{i}$, and simply adding vectors to $\mathbf{h}_{i}$ would draw it away from its valid subspace. To maintain intervened $\mathbf{h}_{i}$ in its valid subspace, it should be rotated along certain direction. The method of precisely changing the encoded number values in language models still remains to be explored.
Figure 17: The success rate of performing a linear intervention on 6 consecutive layers, with a negative $\alpha=-2.0$
We also experimented on negative $\alpha$ values, which will "push" the residual stream towards a smaller encoded number. The results are demonstrated in Figure 17. We can see that the trend of success rate is similar to the trend in Figure 8, further proving that the value of calculation result can be linearly intervened.
@@
Appendix J Directly Calculate with Encoded Number Values